# Sparse Mixture of Experts as Unified Competitive Learning

### v1

@@ 4.5.2 Robustness to the controlling factor
The Unified Competitive Score ($\alpha$) enables the model to adjust its scoring mechanism, either favoring a diverse set of experts per sequence or distributing experts more evenly across tokens. We evaluate the robustness of the controlling factor $\alpha$ on the classification task using the DeepSeekMoE-16B model, with results presented in Table 7. When $\alpha=0.0$, the scoring mechanism aligns with Token Choice, while at $\alpha=1.0$, it follows Expert Choice.
Overall, USMoE demonstrates strong effectiveness within the range of $\alpha\in(0.3,0.7)$, striking a balance between expert diversity and token importance. This range provides an optimal trade-off between enforcing SMoE’s competition policy and enhancing traditional approaches for the task, leading to superior overall performance. Notably, all tested $\alpha$ configurations outperform the Expert Choice approach ($\alpha=1.0$).
Summary and hyper-parameter configuration guidelines. Our experiments demonstrate that USMoE not only excels in pre-training and transfer learning but also achieves superior generalization compared to traditional approaches, all without requiring additional training. Ablation studies in Section 4.5 indicate that USMoE is highly effective within the range of $\alpha\in(0.3,0.7)$. This finding provides a practical guideline for efficiently tuning USMoE through cross-validation.
5 Conclusion
In this research, we reformulate Token Choice and Expert Choice Sparse Mixture of Experts from a competitive learning perspective, highlighting their limitations. Building on this analysis, we introduce Unified Competitive Learning SMoE (USMoE) - a novel and efficient framework that enhances SMoE through a unified competitive learning approach.
We theoretically prove that USMoE achieves superior expert selection compared to traditional methods, effectively improving expert learning capacity while mitigating expert collapse. As a result, USMoE learns more robust expert representations and overcomes the representation collapse issues commonly observed in conventional SMoE training.
Experiments across both training-free and training-based settings (including pre-training and fine-tuning) demonstrate that USMoE enables more efficient and effective training and inference compared to state-of-the-art routing strategies.
Limitations
Ethics Statement
Appendix A Appendix
Supplementary Material for “Sparse Mixture of Experts as Unified Competitive Learning”
This document is organized as follows: Appendix A.1 provides the theoretical proof supporting the discussion in Section 3.2. Following this, Appendix A.3 offers a detailed analysis of the SMoE router and explains why our method outperforms the baseline approaches. Appendix A.2 presents supplementary experimental results, and Appendix A.4 describes the implementation details in full.
A.1 Theoretical Proof for Section  3.2
Let $x\in\mathbb{R}^{N\times d}$ represent an input token embedding and $\mathcal{E}\in\mathbb{R}^{d\times n}$ represent the expert embeddings, where $d$ is the dimension of the SMoE model and $n$ is the number of experts. The similarity score $S$ between the input token and the expert embeddings is computed using a dot product:
$S=x\cdot\mathcal{E}.$ (12)
Since the softmax function is monotonic, the selection of top-$k$ values remains unchanged under softmax transformation. Formally, we have:
By setting $k=1$, the top similarity score for a given token using the Token Choice approach is:
$S^{t}_{j}=\max(S_{ik}),\quad\forall k\in[1,n].$ (14)
Since this operation is performed independently for each input token $i$, we extend this to all tokens $q$ in the batch:
$S^{t}_{j}\leq\max(S_{qk}),\quad\forall k\in[1,n],\quad\forall q\in[1,N].$ (15)
By definition, the Unified Competitive Mechanism approach selects the highest similarity score across all tokens and experts:
$S^{u}_{j}=\max(S_{qk}),\quad\forall k\in[1,n],\quad\forall q\in[1,N].$ (16)
Thus, we establish the first inequality:
$S^{t}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N].$ (17)
Similarly, in the Expert Choice approach, each expert selects the best matching token, leading to:
$S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N].$ (18)
Therefore, we conclude:
$S^{t}_{i}\leq S^{u}_{i},\quad S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N].$ (19)
This completes the proof. $\square$
A.2 Additional Results
We provide a detailed evaluation of three state-of-the-art SMoE models: OLMoE-1B-7B (Table  8), Qwen1.5-MoE-A2.7B (Table 9), and DeepSeekMoE-16B (Table  10). Our results demonstrate the effectiveness of our method across various models and prompts, comparing its performance against baseline approaches such as Token Choice (TC) and Expert Choice (EC).
Category Model Dataset Setting Router TC EC MoEE USMoE
Classification OLMoE-1B-7B Emotion None 24.1 24.5 26.3 25.1 28.2
Prompt 27.6 49.9 49.0 44.5 51.7
Toxic None 51.9 58.9 59.9 51.9 60.6
Prompt 52.3 65.2 61.3 53.4 66.6
Tweet None 47.7 46.8 48.3 48.4 52.1
Prompt 49.5 58.0 58.4 57.2 59.2
Clustering OLMoE-1B-7B Medrxiv None 15.0 17.6 14.8 17.4 21.2
Prompt 15.8 23.9 27.7 22.0 26.3
20Groups None 12.4 11.8 9.2 11.5 12.6
Prompt 16.7 25.7 26.2 24.4 34.6
Pair Classification OLMoE-1B-7B SemEval None 43.6 35.8 31.3 43.6 45.7
Prompt 45.7 46.7 40.9 53.8 54.2
URLCorpus None 47.0 42.4 39.7 47.8 54.5
Prompt 61.4 77.4 76.9 78.2 79.5
Reranking OLMoE-1B-7B Ask None 41.3 41.0 39.0 41.4 42.4
Prompt 43.4 51.9 49.9 50.2 51.8
SciDocs None 45.5 46.3 46.9 50.8 59.9
Prompt 53.6 69.6 73.1 75.1 76.6
StackOver None 25.8 24.8 20.1 26.4 27.0
Prompt 28.1 32.5 30.0 34.3 35.6
STS OLMoE-1B-7B Biosses None 39.3 13.6 7.7 29.7 45.4
Prompt 51.2 61.8 67.6 70.2 74.3
SickR None 50.3 46.3 26.4 53.0 55.1
Prompt 51.9 65.7 37.6 66.1 67.9
STS12 None 40.1 8.6 11.1 37.8 44.6
Prompt 51.3 53.8 37.5 63.6 67.1
STS13 None 40.5 21.1 18.2 43.4 49.8
Prompt 52.5 66.5 40.4 72.7 76.2
STS14 None 29.5 13.4 13.3 31.7 38.7
Prompt 41.1 56.8 33.9 64.2 67.3
STS15 None 30.8 27.8 22.5 33.3 43.3
Prompt 46.4 69.3 38.4 66.4 71.7
STS16 None 46.5 38.9 28.9 45.8 51.7
Prompt 52.4 70.1 49.4 68.3 72.7
STSBen None 42.2 23.4 17.5 44.5 51.1
Prompt 48.6 63.6 48.9 70.7 71.4
Summarization OLMoE-1B-7B Medrxiv None 28.4 20.9 21.1 29.8 30.8
Prompt 25.6 28.9 29.7 30.4 30.9
Table 8: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks with OLMoE-1B-7B models. The best result for each row is highlighted in bold.
Category Model Dataset Setting Router TC EC MoEE USMoE
Classification Qwen1.5-MoE-A2.7B Emotion None 27.2 33.9 30.1 34.3 35.7
Prompt 37.0 48.5 47.4 47.2 49.9
Toxic None 53.0 61.1 21.3 52.9 61.4
Prompt 53.4 64.5 19.8 54.1 66.2
Tweet None 51.1 55.9 25.1 55.9 56.5
Prompt 56.1 61.1 38.6 60.7 61.5
Clustering Qwen1.5-MoE-A2.7B Medrxiv None 15.3 23.3 21.3 23.0 23.4
Prompt 14.2 24.6 19.8 21.8 25.1
20Groups None 12.0 31.5 25.1 27.4 30.0
Prompt 14.4 43.8 38.6 38.4 46.4
Pair Classification Qwen1.5-MoE-A2.7B SemEval None 42.0 38.8 34.7 42.5 44.2
Prompt 47.0 52.4 46.3 52.4 54.1
URLCorpus None 49.8 54.9 52.1 60.6 62.1
Prompt 56.7 68.7 65.8 68.2 72.7
Reranking Qwen1.5-MoE-A2.7B Ask None 43.1 45.8 43.4 47.3 47.5
Prompt 43.3 48.3 49.1 49.5 51.4
SciDocs None 49.6 60.6 55.3 67.0 65.6
Prompt 50.9 60.1 55.8 68.7 73.0
StackOver None 26.2 29.5 26.2 31.1 31.4
Prompt 28.8 31.3 30.2 35.2 36.6
STS Qwen1.5-MoE-A2.7B Biosses None 33.8 32.5 34.7 49.6 52.6
Prompt 55.1 55.8 48.5 68.4 66.2
SickR None 51.0 55.5 40.4 61.0 63.6
Prompt 50.2 59.7 51.1 64.3 66.3
STS12 None 40.2 16.9 18.6 46.3 48.3
Prompt 49.3 25.0 31.8 59.2 61.4
STS13 None 38.1 42.9 44.2 56.7 61.8
Prompt 53.3 57.5 54.6 73.4 75.7
STS14 None 28.1 26.5 25.6 45.4 48.9
Prompt 40.4 38.8 40.7 60.0 62.7
STS15 None 34.8 40.5 38.4 46.1 48.0
Prompt 40.7 52.3 54.2 58.8 62.5
STS16 None 47.6 51.0 48.1 58.1 59.6
Prompt 51.6 64.2 65.1 65.7 68.2
STSBen None 37.0 37.7 34.7 50.9 55.5
Prompt 45.6 47.8 54.5 64.5 67.3
Summarization Qwen1.5-MoE-A2.7B Medrxiv None 28.3 13.4 15.1 31.2 29.7
Prompt 27.0 23.0 21.9 27.3 48.4
Category Model Dataset Setting Router TC EC MoEE USMoE
Classification DeepSeekMoE-16B Emotion None 26.1 27.4 26.5 27.6 27.8
Prompt 37.9 48.3 46.1 46.4 48.8
Toxic None 53.3 60.4 58.1 53.1 60.1
Prompt 53.1 62.4 62.5 53.6 64.2
Tweet None 51.0 51.9 49.5 52.6 52.5
Prompt 54.9 58.4 57.5 58.9 58.9
Clustering DeepSeekMoE-16B Medrxiv None 15.1 23.0 17.3 22.0 25.0
Prompt 17.0 25.7 20.9 24.0 25.8
20Groups None 11.7 13.2 9.7 13.7 18.7
Prompt 18.6 32.3 19.8 33.0 37.9
Pair Classification DeepSeekMoE-16B SemEval None 44.6 40.2 32.6 43.5 46.0
Prompt 48.4 47.2 46.6 51.3 54.6
URLCorpus None 46.4 41.7 41.6 48.6 57.9
Prompt 66.5 72.4 61.1 75.4 75.9
Reranking DeepSeekMoE-16B Ask None 41.7 41.1 40.1 42.3 45.0
Prompt 43.5 43.8 44.7 46.9 49.9
SciDocs None 48.2 50.6 44.7 57.1 61.9
Prompt 58.3 65.6 55.3 72.6 72.9
StackOver None 25.7 24.9 20.4 27.3 30.2
Prompt 29.7 27.6 22.6 32.3 33.6
STS DeepSeekMoE-16B Biosses None 29.5 31.7 27.7 26.8 35.4
Prompt 47.0 40.1 41.5 57.6 55.3
SickR None 50.4 47.4 29.4 53.1 56.7
Prompt 56.0 61.9 38.7 65.8 67.7
STS12 None 44.0 4.3 13.9 45.0 46.9
Prompt 57.8 31.0 28.4 64.0 64.2
STS13 None 36.0 28.4 27.5 41.1 49.1
Prompt 55.3 56.0 41.2 70.9 75.2
STS14 None 25.4 12.0 13.0 28.2 37.3
Prompt 44.9 41.0 31.1 58.6 63.8
STS15 None 34.8 33.9 25.6 38.7 40.7
Prompt 49.7 46.5 33.0 58.5 63.0
STS16 None 44.9 34.4 33.1 46.9 55.7
Prompt 56.7 58.0 44.2 64.5 69.7
STSBen None 36.6 18.3 15.8 42.1 45.6
Prompt 54.9 57.7 39.0 67.8 69.2
Summarization DeepSeekMoE-16B Medrxiv None 24.9 22.0 18.5 24.4 28.4
Prompt 29.1 24.4 25.7 29.2 30.9
A.3 In-depth Analysis
We visualize the router behavior of USMoE in Figure 6 and contrast it with the router behaviors of the Token Choice approach (Figure 4) and the Expert Choice approach (Figure 5). Notably, the router in the OLMoE-1B-7B model exhibits a strong preference for specific experts. For instance, in the Emotion Classification task, Experts 8, 30, and 58 are consistently prioritized in both the Token Choice and Expert Choice approaches. This bias limits the model’s adaptability and effectiveness for downstream tasks. USMoE tackles this challenge by introducing the Unified Competitive Mechanism, which promotes more balanced and diverse expert selections, as illustrated in Figure 6. This enhancement enables USMoE to outperform the baselines on the Emotion Classification task.
@@ A.3 In-depth Analysis
A.4 Implementation Details
For the Without Training experiments, we implement our method based on the publicly available MoEE implementation Li and Zhou (2024) [^1] https://github.com/tianyi-lab/MoE-Embedding. Due to resource constraints, we validate our method and the baselines using 4-bit quantization with a batch size of 128. For the OLMoE-1B-7B model, we conduct experiments on a single H100 GPU, while for the Qwen1.5-MoE-A2.7B and DeepSeekMoE-16B models, we utilize two H100 GPUs.
The base Transformer-XL variant (Chen et al., 2023b) comprises four Transformer decoder layers, each with an input dimension of 256. Each layer includes a self-attention mechanism with eight attention heads, followed by a Feed-forward Neural Network (FFN) that has an inner dimension of 512. The dropout ratio is set at 0.1. We divide the FFN into 16 experts, each with the same dimensions. For the larger variants, we scale the model up to twelve layers.
Our experiments are based on the publicly available SMoE-Dropout implementation (Chen et al., 2023b) [^2] https://github.com/VITA-Group/Random-MoE-as-Dropout. The pre-training experiments were conducted using a single H100 GPU, while the fine-tuning experiments were performed on a single A100 GPU. It is important to note that parallel training on multiple GPUs may produce different results.
A.4.1 Pre-training Experiments
@@
A.4.2 Fine-tuning Experiments

### v2

@@ A.4.2 Training-free additional results
A.4.3 Training from scratch
Transformer-XL(420M) WikiText-103 lm1b
Topkk TC EC USMoE TC EC USMoE
1 31.70 35.52 25.48 58.65 65.43 56.90
2 22.42 23.30 22.06 44.56 43.39 40.53
4 23.57 23.60 22.65 45.52 43.70 40.90
8 24.20 24.37 22.88 46.36 44.22 43.24
[…]
USMoE (Topkk=2) 7.7620 81.5 40.1 88.5 87.8
       (Topkk=1.5) 6.6753 83.8 39.6 88.3 83.0
TC (Topkk=2) SMoE 7.7620 77.1 35.1 84.4 69.2
SMoE-DR 78.6 34.4 83.5 66.7
XMoE 76.7 35.3 83.3 67.4
StableMoE 77.7 34.3 83.9 60.8
EC (Topkk=2) 7.7620 81.5 39.3 88.0 75.6
Training from scratch.  To assess the effectiveness of our method, we compare USMoE with the Token Choice approaches, including SMoE [27], SMoE-Dropout (abbreviated as "SMoE-DR"), XMoE [10], and StableMoE [13], as well as the Expert Choice approach [58] for pre-training and fine-tuning tasks. We follow the approach of [7] and use a base Transformer-XL [14] with four decoder layers. We train both base and large-scale versions of Transformer-XL on four datasets (Enwik8, Text8, Wikitext-103, and One Billion Words) for 100k iterations, following the implementation in [7]. Then we fine-tune the pre-trained weights for text classification tasks, including SST-2 [49], SST-5 [49], IMDB [36], and BANKING77 [4]. More implementation details and additional results are provided in the Appendix A.
Fine-tuning.  We report the results of the fine-tuning experiment on the SST-2, SST-5, IMDB, and BANKING77 datasets in Table  13, using Transformer-XL pre-trained on enwik8. Overall, USMoE consistently achieves higher accuracy compared to other baselines across all datasets. The results demonstrate that our method is not only effective for pre-training tasks but also performs effectively on existing pre-trained models.
A.5 In-depth Analysis
We visualize the router behavior of USMoE in Figure 11 and contrast it with the router behaviors of the Token Choice approach (Figure 9) and the Expert Choice approach (Figure 10). Notably, the router in the OLMoE-1B-7B model exhibits a strong preference for specific experts. For instance, in the Emotion Classification task, Experts 8, 30, and 58 are consistently prioritized in both the Token Choice and Expert Choice approaches. This bias limits the model’s adaptability and effectiveness for downstream tasks. USMoE tackles this challenge by introducing the Unified Mechanism, which promotes more balanced and diverse expert selections, as illustrated in Figure 11. This enhancement enables USMoE to outperform the baselines on the Emotion Classification task.
@@ A.5 In-depth Analysis
A.6 Implementation Details
For the Without Training experiments, we implement our method based on the publicly available MoEE implementation [34] [^1] https://github.com/tianyi-lab/MoE-Embedding. Due to resource constraints, we validate our method and the baselines using 4-bit quantization with a batch size of 128. For the OLMoE-1B-7B model, we conduct experiments on a single H100 GPU, while for the Qwen1.5-MoE-A2.7B and DeepSeekMoE-16B models, we utilize two H100 GPUs.
The base Transformer-XL variant [7] comprises four Transformer decoder layers, each with an input dimension of 256. Each layer includes a self-attention mechanism with eight attention heads, followed by a Feed-forward Neural Network (FFN) that has an inner dimension of 512. The dropout ratio is set at 0.1. We divide the FFN into 16 experts, each with the same dimensions. For the larger variants, we scale the model up to twelve layers.
Our experiments are based on the publicly available SMoE-Dropout implementation [7] [^2] https://github.com/VITA-Group/Random-MoE-as-Dropout. The pre-training experiments were conducted using a single H100 GPU, while the fine-tuning experiments were performed on a single A100 GPU. It is important to note that parallel training on multiple GPUs may produce different results.
A.6.1 Pre-training Experiments
Transformer-XL(20M) Enwik8 Text8 WikiText-103 lm1b
USMoE (Topkk=2) 1.18 1.20 29.20 56.90
       (Topkk=1.5) 1.19 1.28 30.67 57.55
TC (Topkk=2) SMoE 1.20 1.29 30.16 58.00
SMoE-DR 1.56 1.56 58.37 93.17
XMoE 1.21 1.28 30.34 58.33
StableMoE 1.20 1.28 29.97 58.25
EC (Topkk=2) 1.18 1.24 29.83 58.60
@@ A.6.1 Pre-training Experiments
A.6.2 Fine-tuning Experiments